OpenStudy (anonymous):

Simplify: 3 (2x+5)-3x+1 A. 3x+16 B. 3x+6 C. -3x+18 D. 19

OpenStudy (anonymous):

b

OpenStudy (anonymous):

6x+15-3x+1 = 3x+16

OpenStudy (anonymous):

wait so its A ?

OpenStudy (anonymous):

Solve for y: -2x-y= -3 A. y=2x+3 B. y= 2x-3 C. y= -2x-3 D. y= -2x+3

OpenStudy (anonymous):

-2x - y = -3 -y = -3 +2x y = -2x + 3 so, D

OpenStudy (anonymous):

yup A

OpenStudy (anonymous):

thankssss

OpenStudy (anonymous):

NP :)